OpenStudy (anonymous):

what is the length of the line segment between a(-10,8) and b(-10,-3)

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

can you do some thing like less complicated I didnt learn some things yet

hartnn (hartnn):

okk...since this one is simpler, because we have SAME x co-ordinates, the length of segment is just difference in y co-ordinates \(d= 8- (-3) =...?\)